# Vacuum cosmological spacetimes without CMC Cauchy surfaces

### v1

@@ Abstract
In this article, we extend a construction of [6] to obtain a large class of vacuum cosmological spacetimes that do not contain any CMC Cauchy surfaces. The allowed spatial topologies for these examples are of the form $M\#M$, where $M$ is any closed, oriented, irreducible $3$-manifold which is not spherical. This complements the recent results of [19], where, instead of initial data methods, global spacetime gluing arguments were used. The study of such examples is sure to yield insight into Bartnik’s cosmological splitting conjecture [1].
@@ 1 Introduction
Let $(M,g)$ be a cosmological spacetime of dimension $\dim M\geq 3$ satisfying the strong energy condition, i.e. $\mathrm{Ric}(v,v)\geq 0$ for all $v\in TM$ timelike. Suppose $(M,g)$ satisfies the generic condition, i.e. for each inextendible causal geodesic $\gamma$ there exists a parameter $t$ for which $\gamma$ is defined such that the tidal force operator $T_{\gamma(t)}M\to T_{\gamma(t)}M$ defined by $v\mapsto R(v,\gamma^{\prime}(t))\gamma^{\prime}(t)$ is not the zero operator. Then there are incomplete causal geodesics in $M$.
Clearly, this result is wrong without the assumption of genericity. Indeed, let $(S,h)$ be any compact Riemannian manifold of dimension $\dim S\geq 2$ with $\mathrm{Ric}_{h}\geq 0$, then the product spacetime (×$(\mathbb{R}\times S,-dt^{2}+h)$ is cosmological, satisfies the strong energy condition, and is geodesically complete. In this example, genericity is violated along the vertical timelike lines $t\mapsto(t,x_{0})$ for any $x_{0}\in S$.
As is often the case in global geometric results which involve curvature assumptions, one may now wonder whether the cosmological Hawking–Penrose singularity theorem is rigid in its assumption of genericity, i.e. whether every causally geodesically complete cosmological spacetime satisfying the strong energy condition is isometric to a product spacetime of the form (×$(\mathbb{R}\times S,-dt^{2}+h)$, with $(S,h)$ a compact Riemannian manifold of nonnegative Ricci curvature. A slightly stronger result (assuming only timelike geodesic completeness) was conjectured by Bartnik [1] and is one of the most significant open problems in spacetime geometry and mathematical General Relativity to date:
@@
Suppose $(M,g)$ is a cosmological spacetime satisfying the strong energy condition. If $(M,g)$ is timelike geodesically complete, then it is isometric to a product spacetime (×$(\mathbb{R}\times S,-dt^{2}+h)$, where $(S,h)$ is a compact Riemannian manifold of nonnegative Ricci curvature.
@@
Our contribution in this paper is to extend the example of [6] by generalizing $T^{3}$ to an arbitrary compact, oriented, irreducible, non-spherical $3$-manifold $M$. Due to the topological nature of the argument, this generalizes the case of $3$ space dimensions and symmetric sides in the gluing, but not the more general situations considered in [19].
@@ 2 The main result
Let $M$ be a closed oriented 3-manifold. $M$ is called prime if it cannot be written as a nontrivial connected sum of two 3-manifolds. $M$ is called irreducible if every embedded 2-sphere bounds a 3-ball. Clearly, $M$ is prime whenever its irreducible. Conversely, if $M$ is prime, then either $M$ is irreducible or it is topologically $\mathbb{S}^{1}\times\mathbb{S}^{2}$ [16, Lem. 3.13]. Suppose $M$ is irreducible. Either $M$ has finite fundamental group or not. In the former case, $M$ is topologically a spherical space by the positive resolution of the elliptization conjecture. [^4] By a spherical space, we mean any quotient of the 3-sphere $\mathbb{S}^{3}/\Gamma$ by a discrete subgroup $\Gamma\subset\text{SO}(4)$ acting freely on $\mathbb{S}^{3}$. Our main theorem concerns the latter case.
@@
For any closed oriented irreducible $3$-manifold $M$ which is not topologically a spherical space, there is a vacuum initial data set on the connected sum $M\#M$ whose maximal globally hyperbolic development does not contain any CMC Cauchy surfaces.
@@
More generally, Theorem 2.1 remains true if $M$ is a closed oriented $3$-manifold whose prime decomposition $M=M_{1}\#\dotsb\#M_{n}$ contains some factor $M_{k}$ that is irreducible and not a spherical space.
@@ 3 The proof
A vacuum initial data set is a triple $(M,g,K)$, where $M$ is a 3-manifold, $g$ is a Riemannian metric on $M$, and $K$ is a symmetric $(0,2)$ tensor on $M$ such that $g$ and $K$ satisfy the vacuum Einstein constraint equations, [^5] We do not consider the cosmological constant in this paper, i.e., $\Lambda=0$.
@@
where $R$ is the scalar curvature of $g$ and $H=\text{tr}_{g}K$ will be referred to as the mean curvature. If $H$ is constant, then $(M,g,K)$ is called a constant mean curvature (CMC) vacuum initial data set.
A pair $(N,Y)$ is called a Killing initial data (KID) for an initial data set $(M,g,K)$ if $N$ is a function on $M$ and $Y$ is a vector field on $M$ such that
@@
where $\mathcal{L}_{Y}K$ is the Lie derivative of $K$ with respect to $Y$ (see [2, Eqs. (5.3) & (5.4)]). If $(\overline{M},\overline{g})$ is the maximal globally hyperbolic vacuum spacetime development of $(M,g,K)$, then KIDs are in one-to-one correspondence with Killing vector fields of $(\overline{M},\overline{g})$ (see [20, 21]).
@@
Suppose $(M,g,K)$ is a CMC vacuum initial data set and $M$ is closed. If $(N,Y)$ is a KID, then $N=0$ and $Y$ is a Killing vector field.
@@
It suffices to show $N=0$. We claim that
@@
Once this is shown, multiplying this equation by $N$ and integrating by parts shows that $N=0$ since $M$ is closed; the left hand side is manifestly nonpositive and the right hand side is manifestly nonnegative.
@@
For any closed 3-manifold $M$, there is a vacuum initial data set $(M,g,K)$ with no nontrivial KIDS.
@@
By [2, Thm. 2.1], there is a Riemannian metric $\hat{g}$ on $M$ such that $\hat{g}$ has no nontrivial local conformal Killing vector fields. Let $\hat{\sigma}$ be a nontrivial traceless and divergence-free tensor with respect to $\hat{g}$. Such tensors exist by general results in [7] (in fact, the space of TT tensors is infinite-dimensional; see Remark 9.9 in [7]). Hence, the conformal method applies [17, Sect. 4], and so there is a vacuum initial data set $(M,g,K)$ such that $g$ is a conformal deformation of $\hat{g}$ and $H=\text{tr}_{g}K$ is constant. If $(N,Y)$ is a KID for $(M,g,K)$, then the lemma implies that $N=0$ and $Y$ is a Killing vector field for $(M,g)$, but this contradicts the fact that $\hat{g}$ has no nontrivial local conformal Killing vector fields. ∎
@@
Proof of Theorem 2.1. By the previous proposition, there is a vacuum CMC initial data set $(M,g,K)$ with no nontrivial KIDs. As shown in [6], the lack of nontrivial KIDs allows one to perform a gluing procedure on $(M_{1},g,K)$ and $(M_{2},g,-K)$ with $M_{i}\approx M$ producing a vacuum initial data set on $M\#M$ such that the metric is symmetric under a reflection across the middle of the connecting neck, while the second fundamental form changes sign under this reflection. Due to this symmetry at the level of the initial data, the resulting maximal globally hyperbolic vacuum spacetime development $(\overline{M},\overline{g})$ admits a time-inverting isometry.
@@
Seeking a contradiction, suppose $(\overline{M},\overline{g})$ contains a CMC Cauchy surface. Then, as in [1] (see also [19]), using this and the time-inverting isometry produces barriers for the existence of maximal Cauchy surface $S$ [1, 12]. The constraint equations imply that $S$ has nonnegative scalar curvature. However, its fundamental group is $\pi_{1}(S)=\pi_{1}(M\#M)=\pi_{1}(M)\ast\pi_{1}(M)$. Since $M$ is irreducible and not a spherical space, it has infinite fundamental group by the positive resolution of the elliptization conjecture [22, 23, 24]. It follows by the positive resolution of the surface subgroup conjecture [18] that there is a subgroup of $\pi_{1}(S)$ which is isomorphic to the fundamental group of a surface of genus $g\geq 1$. By [25, Thm. 5.2], it follows that $S$ is a flat closed oriented 3-manifold, but these have been classified and all of them are prime. To see this, recognize that a flat closed 3-manifold is a Seifert fibre space (see [26, Thm. 4.3]). But the only Seifert fibre space which is a nontrivial connected sum is $\mathbb{R}\mathbb{P}^{3}\#\mathbb{R}\mathbb{P}^{3}$ (see [13, Prop. 1.12]). But $\mathbb{R}\mathbb{P}^{3}\#\mathbb{R}\mathbb{P}^{3}$ is double covered by $\mathbb{S}^{1}\times\mathbb{S}^{2}$ which does not carry a flat metric [27, Thm. 3.3.1]. (Alternatively, $S$ being $\mathbb{R}\mathbb{P}^{3}\#\mathbb{R}\mathbb{P}^{3}$ topologically would contradict the fact that $M$ is not a spherical space.) ∎
@@ 4 Outlook
In this article, we generalized the example of a vacuum cosmological spacetime without CMC Cauchy surfaces obtained in [6] to initial data on the connected sum $M\#M$ for any closed oriented, irreducible, non-spherical $3$-manifolds $M$. Our topological argument makes use of the positive resolutions of the elliptization conjecture [22, 23, 24] and the surface subgroup conjecture [18]. Along with examples obtained in [19], we have a plethora of cosmological spacetimes without CMC Cauchy surfaces with varying topologies. But all examples constructed so far have Cauchy surfaces on nonprime 3-manifolds, so it would be interesting to find examples of cosmological spacetimes without CMC Cauchy surfaces whose Cauchy surfaces are prime 3-manifolds.

### v2

@@ Abstract
In this article, we extend a construction of [6] to obtain a large class of vacuum cosmological spacetimes that do not contain any CMC Cauchy surfaces. The allowed spatial topologies for these examples are of the form $M\#M$, where $M$ is any closed, connected, oriented, irreducible $3$-manifold which is not spherical. This complements the recent results of [19], where, instead of initial data methods, global spacetime gluing arguments were used. The study of such examples is sure to yield insight into Bartnik’s cosmological splitting conjecture [1].
@@ 1 Introduction
Let $(\tilde{M},\tilde{g})$ be a cosmological spacetime of dimension $\dim\tilde{M}\geq 3$ satisfying the strong energy condition, i.e., $\mathrm{Ric}(v,v)\geq 0$ for all $v\in T\tilde{M}$ timelike. Suppose $(\tilde{M},\tilde{g})$ satisfies the generic condition, i.e., for each inextendible causal geodesic $\gamma$ there exists a parameter $t$ for which $\gamma$ is defined such that the tidal force operator $T_{\gamma(t)}\tilde{M}\to T_{\gamma(t)}\tilde{M}$ defined by $v\mapsto R(v,\gamma^{\prime}(t))\gamma^{\prime}(t)$ is not the zero operator. Then there are incomplete causal geodesics in $\tilde{M}$.
Clearly, this result is wrong without the assumption of genericity. Indeed, let $(S,g)$ be any compact Riemannian manifold of dimension $\dim S\geq 2$ with $\mathrm{Ric}_{g}\geq 0$, then the product spacetime (×$(\mathbb{R}\times S,-dt^{2}+g)$ is cosmological, satisfies the strong energy condition, and is geodesically complete. In this example, genericity is violated along the vertical timelike lines $t\mapsto(t,x_{0})$ for any $x_{0}\in S$.
As is often the case in global geometric results which involve curvature assumptions, one may now wonder whether the cosmological Hawking–Penrose singularity theorem is rigid in its assumption of genericity, i.e., whether every causally geodesically complete cosmological spacetime satisfying the strong energy condition is isometric to a product spacetime of the form (×$(\mathbb{R}\times S,-dt^{2}+g)$, with $(S,g)$ a compact Riemannian manifold of nonnegative Ricci curvature. A slightly stronger result (assuming only timelike geodesic completeness) was conjectured by Bartnik [1] and is one of the most significant open problems in spacetime geometry and mathematical General Relativity to date:
@@
Suppose $(\tilde{M},\tilde{g})$ is a cosmological spacetime of dimension $\dim\tilde{M}\geq 2$ satisfying the strong energy condition. If $(\tilde{M},\tilde{g})$ is timelike geodesically complete, then it is isometric to a product spacetime (×$(\mathbb{R}\times S,-dt^{2}+g)$, where $(S,g)$ is a compact Riemannian manifold of nonnegative Ricci curvature.
@@
Our contribution in this paper is to extend the example of [6] by generalizing $T^{3}$ to an arbitrary compact, connected, oriented, irreducible, non-spherical $3$-manifold $M$. Due to the topological nature of the argument, this generalizes the case of $3$ space dimensions and symmetric sides in the gluing, but not the more general situations considered in [19].
@@ 2 The main result
Let $M$ be a closed, connected, oriented 3-manifold. $M$ is called prime if it cannot be written as a nontrivial connected sum of two 3-manifolds. $M$ is called irreducible if every embedded 2-sphere bounds a 3-ball. Clearly, $M$ is prime whenever its irreducible. Conversely, if $M$ is prime, then either $M$ is irreducible or it is topologically $\mathbb{S}^{1}\times\mathbb{S}^{2}$ [16, Lem. 3.13]. Suppose $M$ is irreducible. Either $M$ has finite fundamental group or not. In the former case, $M$ is topologically a spherical space by the positive resolution of the elliptization conjecture. [^4] By a spherical space, we mean any quotient of the 3-sphere $\mathbb{S}^{3}/\Gamma$ by a discrete subgroup $\Gamma\subset\text{SO}(4)$ acting freely on $\mathbb{S}^{3}$. Our main theorem concerns the latter case.
@@
For any closed, connected, oriented, irreducible $3$-manifold $M$ which is not topologically a spherical space, there is a vacuum initial data set on the connected sum $M\#M$ whose maximal globally hyperbolic development does not contain any CMC Cauchy surfaces.
@@
More generally, Theorem 2.1 remains true if $M$ is a closed, connected, oriented $3$-manifold whose prime decomposition $M=M_{1}\#\dotsb\#M_{n}$ contains some factor $M_{k}$ that is irreducible and not a spherical space.
@@ 3 The proof
A vacuum initial data set is a triple $(S,g,K)$, where $S$ is a smooth manifold of dimension $\dim S\geq 2$, $g$ is a Riemannian metric on $S$, and $K$ is a symmetric $(0,2)$ tensor on $S$ such that $g$ and $K$ satisfy the vacuum Einstein constraint equations, [^5] We do not consider the cosmological constant in this paper, i.e., $\Lambda=0$.
@@
where $R$ is the scalar curvature of $g$ and $H=\text{tr}_{g}K$ will be referred to as the mean curvature. If $H$ is constant, then $(S,g,K)$ is called a constant mean curvature (CMC) vacuum initial data set.
A pair $(N,Y)$ is called a Killing initial data (KID) for an initial data set $(S,g,K)$ if $N$ is a function on $S$ and $Y$ is a vector field on $S$ such that
@@
where $\mathcal{L}_{Y}K$ is the Lie derivative of $K$ with respect to $Y$ (see [2, Eqs. (5.3) & (5.4)]). If $(\tilde{M},\tilde{g})$ is the maximal globally hyperbolic vacuum spacetime development of $(S,g,K)$, then KIDs are in one-to-one correspondence with Killing vector fields of $(\tilde{M},\tilde{g})$ (see [20, 21]).
@@
Let $(S,g,K)$ be a CMC vacuum initial data set with $S$ closed, connected and $\dim S\geq 2$. If $(N,Y)$ is a KID for $(S,g,K)$, then $Y$ is a Killing vector field. Moreover, $N$ is constant on $S$, and $N\neq 0$ implies $K=0$.
@@
It suffices to show the claims about $N$, since either of the options imply that $Y$ is a Killing vector field. We claim that
@@
Once this is shown, multiplying this equation by $N$ and integrating by parts shows that $\nabla N=0$ since $S$ is closed, hence $N$ is constant; the left hand side is manifestly nonpositive and the right hand side is manifestly nonnegative. Since the constancy of $N$ trivially implies $\Delta N=0$, we either have that $N=0$, or else $K=0$.
@@
For any closed, connected smooth manifold $S$ of dimension $\dim S\geq 3$, there is a vacuum initial data set $(S,g,K)$ with no nontrivial KIDs.
@@
By [2, Thm. 2.1, Thm. 7.4], there is a Riemannian metric $\hat{g}$ on $S$ such that $\hat{g}$ has no nontrivial local conformal Killing vector fields. Let $\hat{\sigma}$ be a nontrivial traceless and divergence-free tensor with respect to $\hat{g}$. Such tensors exist by general results in [7] (in fact, the space of TT tensors is infinite-dimensional; see Remark 9.9 in [7]). Hence, the conformal method applies [17, Sect. 4], and so there is a vacuum initial data set $(S,g,K)$ such that $g$ is a conformal deformation of $\hat{g}$ and $H=\text{tr}_{g}K$ is constant, which we may specify to be nonzero. If $(N,Y)$ is a KID for $(S,g,K)$, then Lemma 3.1 implies that $Y$ is a Killing vector field for $(S,g)$, as well as either $N=0$ or else $N$ is a nonzero constant and $K=0$. Since $\hat{g}$ has no nontrivial local conformal Killing vector fields, it follows that $Y=0$. Moreover, $N\neq 0$ and thus $K=0$ is ruled out since we specified $H\neq 0$ in the application of the conformal method. Thus, $N=0$ and $Y=0$, as claimed. ∎
Let us now come to the proof of Theorem 2.1 and recall that, in the following, $M$ denotes a closed, connected, $3$-dimensional, oriented, irreducible smooth manifold which is topologically not a spherical space, i.e., not diffeomorphic to a homogeneous space of the form $\mathbb{S}^{3}/\Gamma$, where $\Gamma\subset SO(4)$ is a discrete subgroup acting freely on $\mathbb{S}^{3}$.
@@
Proof of Theorem 2.1. By Proposition 3.2, there is a vacuum CMC initial data set $(M,g,K)$ with no nontrivial KIDs. As shown in [6], the lack of nontrivial KIDs allows one to perform a gluing procedure on $(M_{1},g,K)$ and $(M_{2},g,-K)$ with $M_{i}\approx M$ producing a vacuum initial data set on $M\#M$ such that the metric is symmetric under a reflection across the middle of the connecting neck, while the second fundamental form changes sign under this reflection. Due to this symmetry at the level of the initial data, the resulting maximal globally hyperbolic vacuum spacetime development $(\tilde{M},\tilde{g})$ admits a time-inverting isometry.
@@
Seeking a contradiction, suppose $(\tilde{M},\tilde{g})$ contains a CMC Cauchy surface. Then, as in [1] (see also [19]), using this and the time-inverting isometry produces barriers for the existence of a maximal Cauchy surface $C$ [1, 12]. The constraint equations imply that $C$ has nonnegative scalar curvature. However, its fundamental group is $\pi_{1}(C)=\pi_{1}(M\#M)=\pi_{1}(M)\ast\pi_{1}(M)$. Since $M$ is irreducible and not a spherical space, it has infinite fundamental group by the positive resolution of the elliptization conjecture [22, 23, 24]. It follows by the positive resolution of the surface subgroup conjecture [18] that there is a subgroup of $\pi_{1}(C)$ which is isomorphic to the fundamental group of a surface of genus $\geq 1$. By [25, Thm. 5.2], it follows that $C$ is a flat, closed, connected, oriented 3-manifold, but these have been classified and all of them are prime. To see this, recognize that a flat, closed, connected, oriented 3-manifold is a Seifert fibre space (see [26, Thm. 4.3]). But the only Seifert fibre space which is a nontrivial connected sum is $\mathbb{R}\mathbb{P}^{3}\#\mathbb{R}\mathbb{P}^{3}$ (see [13, Prop. 1.12]). But $\mathbb{R}\mathbb{P}^{3}\#\mathbb{R}\mathbb{P}^{3}$ is double covered by $\mathbb{S}^{1}\times\mathbb{S}^{2}$ which does not carry a flat metric [27, Thm. 3.3.1]. Alternatively, $C$ being $\mathbb{R}\mathbb{P}^{3}\#\mathbb{R}\mathbb{P}^{3}$ topologically would contradict the fact that $M$ is not a spherical space. ∎
@@ 4 Outlook
In this article, we generalized the example of a vacuum cosmological spacetime without CMC Cauchy surfaces obtained in [6] to initial data on the connected sum $M\#M$ for any closed, connected, oriented, irreducible, non-spherical $3$-manifold $M$. Our topological argument makes use of the positive resolutions of the elliptization conjecture [22, 23, 24] and the surface subgroup conjecture [18]. Along with examples obtained in [19], we have a plethora of cosmological spacetimes without CMC Cauchy surfaces with varying topologies. But all examples constructed so far have Cauchy surfaces on nonprime 3-manifolds, so it would be interesting to find examples of cosmological spacetimes without CMC Cauchy surfaces whose Cauchy surfaces are prime 3-manifolds.